OpenStudy (anonymous):

A combined circuit has two resistors in parallel (10.0 ohms and 14.0 ohms) and another in series (5.0 ohms). If the power source is 9.0 volts, what is the current? A. 1.6 amps B. 5.8 amps C. 0.83 amps D. 1.3 amps

OpenStudy (anonymous):

Odd. My reply vanished. Parallel 1/R' = 1/10+1/14=.171 R' = 5.83 The, series: R = 5.83 + 5 = 10.83 current I = V/R = 9/10.83 = 0.83 amps